OpenStudy (anonymous):

HOW DO I TYPE THIS IN ACTUAL MATH VIEW: the square root of x to the fourth power: (x^4)^1/2

OpenStudy (anonymous):

\[\sqrt{x^4}\] like that?

OpenStudy (anonymous):

same thing as \(x^2\)

OpenStudy (anonymous):

Yes like that ! I need to write (x^4)^1/2

OpenStudy (anonymous):

you should just write \(x^2\) since \(\large (x^4)^{\frac{1}{2}}=x^{4\times \frac{1}{2}}=x^2\)

OpenStudy (anonymous):

How about this one ( In the same way you wrote it above) and yes I know they all end up equaling x to the second power. (x^8)^1/4

OpenStudy (anonymous):

i am not sure what you are asking

OpenStudy (anonymous):

just like you did above, I need to do this one now (x^8)^1/4

OpenStudy (anonymous):

\[(x^{8})^{\frac{1}{4}}=x^{8\times \frac{1}{4}}=x^{\frac{8}{4}}=x^2\]

OpenStudy (anonymous):

and this last one please ! (x^16)^1/8

OpenStudy (anonymous):

thankyouu ! Thats the last one I need to convert, please help !

OpenStudy (anonymous):

\[\large (x^{16})^{\frac{1}{8}}=x^{16\times \frac{1}{8}}=x^{\frac{16}{8}}=x^2\]